OpenStudy (anonymous):

the sum of two consecutive even integers is 134 what are the integers?

OpenStudy (xguardians):

x + (x+2) = 134 2x + 2 = 134 2x = 132 x = 66 66, 68

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm x+(x+2)=134 }\) \(\large\color{midnightblue}{ \rm x+x+2=134 }\) \(\large\color{midnightblue}{ \rm 2x+2=134 }\) \(\large\color{midnightblue}{ \rm 2x=132 }\) \(\large\color{midnightblue}{ \rm x=66 }\) \(\large\color{midnightblue}{ \rm x=66~~~~~~~~~~(x+2)=68 }\)

OpenStudy (xguardians):

Why did you write exactly what I did lol.